Grade 11Physical Chemistry

Iodine molecule dissociates into atoms after absorbing light of 4500 Å. If one quantum of radiation is absorbed by each molecule, calculate the kinetic energy of iodine atoms.
(Bond energy of I2 = 240 kJ mol-1)

Simran Bhatia
12 Years agoGrade 11

1 Answer

Aditi Chauhan
12 Years ago
Bond energy of I2 240kJ mol-1 = 240 * 103 J mol-1
= 240 *103/6.023 *1023 J molecule-1
= 3.984 *10-19 J molecule-1
Energy absorbed = hc/λ = 6.626 *10-34 Js *3 *108 ms-1/4500 *10-10 m
= 4.417 * 10-19 J
Kinetic energy = Absorbed energy – Bond energy
∴ Kinetic energy = 4.417 *10-19 – 3.984 *10-19 J
= 4.33 * 10-20 J
∴ Kinetic energy of each atom of iodine
= 4.33 *10-20/2 = 2.165 *10-20